OpenStudy (eric_d):

Sketch graph of y=x^2+3/2 @rational

OpenStudy (rational):

what do you know about the given function ?

OpenStudy (eric_d):

f(0)= 3/2

OpenStudy (eric_d):

It's a quadratic function

OpenStudy (rational):

yes thats the y intercept - but even before that, notice that the given function is a quadratic. So what happens at the extreme values of x ?

OpenStudy (eric_d):

I don't know

OpenStudy (harsha19111999):

Hey take a value of x and see what the value of y is. Then you get some points

OpenStudy (rational):

Since the leading coefficient is 1, which is positive, at extreme values of x, the graph goes UP

OpenStudy (eric_d):

Okay, I don't hv to find the value of x

OpenStudy (harsha19111999):

For example, if x=0, y =3/2 and if x=1, y=1+3/2=5/2 and if x=2, y=4+3/2=11/2

OpenStudy (harsha19111999):

With these points (0,3/2),(1,5/2),(2,11/2) draw a graph

OpenStudy (rational):



OpenStudy (rational):

and you know that the y intercept = 3/2

OpenStudy (eric_d):